OpenStudy (goformit100):

Explain why 3 × 5 × 7 + 7 is a composite number.

OpenStudy (goformit100):

@genius12

OpenStudy (anonymous):

because it has a factor of 7

OpenStudy (goformit100):

But how it has the factor, Sir please show.

OpenStudy (anonymous):

\(3\times 5\times 7+7=(3\times 5+1)\times 7\) by the distributive law

OpenStudy (anonymous):

of course you could always compute the number and see that it is composite

OpenStudy (goformit100):

OMG! i got it now :) how foolish I was, didn't use or thought about distributive law

OpenStudy (goformit100):

Anyways, Thank you

OpenStudy (anonymous):

yw